Solve the following equations. 
 9x/(7 – 6x) = 15 
Solve the following equations. 
 9x/(7 – 6x) = 15 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

9x/(7–6x)=15
⇒9x=15(7-6x)
⇒9x=105-90x
⇒9x+90x=105
⇒99x=105
⇒x=105/99
=35/33

Thanks